An Urban council has installed 2000 lamps with mercury bulbs in the   streets of town area. The lifetimes of these bulbs are normally distributed with a mean of 1200 burning hours and having a standard deviation of 200 hours. 

(i) How many of these bulbs can be expected to fail.

 (a) In the first 800 burning hours? 



1
Expert's answer
2020-12-29T12:50:25-0500

Let random variable X represents life time of the bulb


X~N("\mu" , "\sigma" 2) =N (1200 , 2002) then Z = (X - "\mu" )/"\sigma" = (X -1200 )/200 ~N (0 , 1)

P(X"\leq"  800) = P((X -1200 )/200"\leq" (800 -1200 )/200)= P (Z"\leq" -2 ) = 0.0228

Number of bulbs expected to fail in the first 800 burning hours

= P(X"\leq" 800)*(total number of lamps)

= 0.0228 (2000) = 45.6

Answer : Number of bulbs expected to fail in the first 800 burning hours is approximately equal to 46.